OpenStudy (anonymous):

1. What is the vertex of the graph? (1 point)(2, 2) (1, 3) (3, 1) (0, 2) 2. Is the vertex of the graph in problem 1 a maximum or a minimum? (1 point)maximum minimum neither both Note: The question below was entered in error. You will receive credit for ANY answer. 3. A ball is thrown into the air with an initial upward velocity of 60 ft/s. Its height (h) in feet after t seconds is given by the function h = –16t² + 60t + 6. After about how many seconds will the ball hit the ground? (1 point)3 4 5 6 4. A ball is thrown into the air with an initial upward veloci

OpenStudy (anonymous):

1. What is the vertex of the graph? (1 point)(2, 2) (1, 3) (3, 1) (0, 2) 2. Is the vertex of the graph in problem 1 a maximum or a minimum? (1 point)maximum minimum neither both Note: The question below was entered in error. You will receive credit for ANY answer. 3. A ball is thrown into the air with an initial upward velocity of 60 ft/s. Its height (h) in feet after t seconds is given by the function h = –16t² + 60t + 6. After about how many seconds will the ball hit the ground? (1 point)3 4 5 6 4. A ball is thrown into the air with an initial upward velocity of 60 ft/s. Its height (h) in feet after t seconds is given by the function h = –16t² + 60t + 6. What will the height be at t = 3 seconds? (1 point)35 feet 40 feet 42 feet 45 feet 5. Solve x² – 121 = 0 (1 point)0 –11 11 11, –11 6. Solve by factoring. m² + 8m + 7 = 0 (1 point)8, 7 –7, 1 –7, –1 7, 1 7. Solve by factoring. n² + 2n – 24 = 0 (1 point)–12, 2 –2, 12 –6, 4 –4, 6 8. One more rectangular-shaped piece of metal siding needs to be cut to cover the exterior of a pole barn. The area of the piece is 30 ft². The length is 1 less than 3 times the width. How wide should the metal piece be? Round to the nearest hundredth of a foot. (1 point)3.33 ft 4.3 ft 1 ft 30 ft

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!